building w 12 units consists of 1bedrooms costing 360 and 2 bedrooms costing 450 respectively. when all units are full total is 4950. how many 2bedroom units?
Answer by ewatrrr(10682)


Hi
12 units
Let x and (12-x) represent 2 and 1 bedroom apt respectively
Question states***
$450x + $360(12-x) = $4950
Solving for x
90x = 4950 - 360*12
90x = 630 x = 7, number of 2 BR. 5 the number of 1 BR. (12-7)
CHECKING our Answer***
7*$450 + 5*$360 = $4950